OpenStudy (anonymous):

What is the slope of the line 5x - 2y = 10?

OpenStudy (anonymous):

-2 y = -5x + 10 y = 5/2x - 5 m = 5/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem!